OpenStudy (anonymous):

Please help! Will fan & medal!! A model rocket is launched from the ground with an initial velocity of 288 ft/sec. How long will it take the rocket to reach its maximum height? Show all work in the space provided.

OpenStudy (anonymous):

@dan815

OpenStudy (welshfella):

to do this youy use one of the equations for constant acceleration the acceleration in this case is that due to gravity which = 32 ft s-2 in this case we use -32 bacause it is acting against the motion of the rocket

OpenStudy (welshfella):

use the formula v = u + at At maximum height the velocity v = 0 , a the acceleration = -32 , the initial velocity = 288 ft/sec and t is the required time solve for t

OpenStudy (welshfella):

initial velocity = u

OpenStudy (anonymous):

Is it 9?

OpenStudy (welshfella):

so just plug those values into the formula

OpenStudy (welshfella):

0 = 288 -32t 32t = 288 t = 9 seconds yes

OpenStudy (anonymous):

Thank you! You helped explained it good. Could I ask another question? I'll make sure to close this one.

OpenStudy (welshfella):

ok